# Accounting for multiplicity in machine learning benchmark performance

### v4

@@ 5.1.2 Cassava Leaf Disease
In this example, the highest reported accuracy is unlikely to be the product of multiplicity. For The Cassava Leaf Disease Classification competition, the task was to classify images of cassava leaves. The training set consisted of $21,367$ labelled images of 5 classes, and the test set of approximately $15,000$ images. $3,752$ teams participated (excluding those with accuracy lower than chance).
The three top performances appear as outliers, as seen in Figure 8(a). The 95% CI of $\hat{\theta}^{\,\max}=0.9132$ is $(0.9086,0.9177)$, no other observations are within this interval. By performing simulations without cropping, we get $E(\hat{\tilde{\theta}}^{\,max})=\hat{\theta}^{\,\max}$, and although the upper tail is stretched, its tip does not surpass $\hat{\theta}^{\,max}$.
@@
Figure 8: The competition observations and one realisation.
Several explanations for the outlying performances of the three top teams have been put forward, and the interested reader can take a deep-dive in the comment section of the competition. Since this is not the topic of the present work, we will not engage in the discussion here.
5.2 Simulation of AUCs
@@
We discuss here an example, the SIIM-ISIC Melanoma Classification competition from 2020 because it follows many of the common structures of public challenges. The winning team chose many of the verified strategies from Section 2.
@@
The SIIM-ISIC Melanoma Classification competition consisted of more than $40,000$ images of skin lesions from various sites, split into training ($33,126$), validation ($3,295$) and test set ($7,687$). The data set is highly imbalanced, with less than $2\%$ melanomas, which reflects the reality in many cancer detection situations. A full description can be found in Rotemberg et al. (2021). The size of the test set is smaller than the recommendations of, e.g., Roelofs et al. (2019), but reflects the reality of benchmark datasets and public challenges, as seen, e.g., in the overview of Willemink et al. (2020). We want to emphasise that the shortcomings of this competition are not unique, and that is precisely why it serves as a useful example. Participants were allowed to use additional data for training if the data was shared on the competition web site. The prize money was set to $10,000$ USD for the best-performing method, measured as AUC on the hold-out test set, published in the private leaderboard as seen in Figure 9. Participants could get intermediate results from the validation set, published on a public leaderboard.
Figure 9: Histogram of the AUCs from the kaggle leaderboard, test set.
@@
The team ‘All Data Are Ext’ consisting of three Kaggle Grandmasters won the competition, achieving an AUC of $0.9490$ on the test set, and their method and strategy is explained in Ha et al. (2020). Although we show that the AUC is an optimistic estimate for the melanoma detection state-of-the-art, we do not question their rank in the challenge. Quite the contrary; the research in Section 2 supports their strategies. They had a low rank on the public leaderboard (880th out of 3300). Their AUC decreased from the public to the private leaderboard, but it was still within the confidence interval. They used additional data from previous years’ competitions, and they also did data augmentation. They chose ensemble learning as a strategy to avoid overfitting, a strategy implicitly supported by Talagrand (1996). They also trained on multiple classes, a strategy that Feldman et al. (2019) showed will slow down the overfitting.
@@
In a binary classification problem, the calculation of AUC is based on scores of random variables from the positive and the negative class. The score function plus a threshold is the classifier. The AUC is mathematically equivalent to the probability that a random variable from the positive class will be ranked higher, in terms of score, than a random variable from the positive class (Hanley and McNeil, 1982).
@@
Let $score(\cdot)$ be a function that assigns a score to a random variable. Let $S_{+}=score(W_{+})$, where $W_{+}$ comes from the positive class, and $S_{-}=score(W_{-})$, where $W_{-}$ comes from the negative class. Then $\text{AUC}=P(S_{-}<S_{+})=P(S_{-}-S_{+}<0)$, which is equivalent to the cdf at value 00. Hence to simulate classifier scores according to a given true or expected AUC, $a$, we simply need to specify distributions for the variables $S_{-},S_{+}$ such a way that $P(S_{Y}-S_{X}<0)=a$.
This is particularly simple with Normal distributions, since linear combinations of Normal distributions yields another Normal. Let $S_{+}$ and $S_{-}$ have the distributions
@@
where $\mathcal{N}(\mu,\sigma)$ is the Normal distribution with mean $\mu$ and variance $\sigma^{2}$. We then have that
@@
by linear combination of two Normal distributions, where $\Phi$ is the cdf of the standard Normal distribution. Specifying any three of the four parameters above decides the last one for a chosen AUC = a. Figure 10(a) shows an example where AUC = 0.90.
@@
The uncertainty of AUC can be calculated through bootstrapping, but without access to the predictions for each data point in the test set, the analyst must make some assumptions about the shape of the ROC curve.
Figure 10(a) shows the pdfs of negative and positive scores with an AUC of 0.90. The rug shows an example draw with 3000 images, with the probability of drawing a positive image is 0.017, corresponding to the class balance in the Melanoma challenge. Fig. 10 shows different shapes of ROC curves.
(a) Score pdfs for a classifier with AUC= $0.90$; grey for negatives, black for positives. The rug shows an example draw from these.
(b) ROC curves for a classifier with expected AUC of $0.90$ across 100 realisations of (a).
Figure 10: Example of what an AUC of $0.90$ might look like in practice. The right figure shows $100$ realizations of a ROC curve from a competition with $3000$. Out of the $3000$ images, $52$ are positives, reflecting the proportion of the Kaggle melanoma challenge. The left figure shows the underlying score distributions that generated the curves of Figure 10(b).
However, with high AUC, the shape has less impact, and the analyst’s choices are not suggestive of the overall conclusion.
@@ 5.2.1 Simulation results for uncorrelated classifiers
For simplicity we choose $\mu_{-}=0$ and $\sigma_{-}=\sigma_{+}=1$. We then solve for $\mu_{+}$ numerically. Now we can generate scores from the two resulting normal distributions.
@@
Whereas the distribution of $\hat{\theta}^{\,\max}$ is independent of the class balance as long as the probability of correct prediction is independent of class, this is not true for maximum of $\hat{AUC}$, where the variance increases with class imbalance. We have chosen the same class balance as in the Melanoma challenge, as this reflects real world examples. The 95% CI of a single classifier with observed AUC equal to its true AUC of $0.90$ is (.$(.8558,0.9376)$, and gives a strong indication of what we can expect when $1,000$ teams apply their classifiers on a test set.
Figure 11: Example distribution of simulated observed state of the art with parameters as in Figure 10. This is an analogous figure to Figure 3 above.
Figure 11 shows the simulated distribution of the observed maximum AUC over $10,000$ repetitions of a competition where every classifier has a true AUC of $0.90$. The simulated expected value is $0.9562$, the standard deviation is $0.004459$, and the simulated 95% CI is $(0.9486,0.9662)$.
To give an unbiased estimate of the SOTA performance for the Melanoma challenge, we follow the same procedure as in Section 5.1: The competition observations are cropped in such a manner that their realisations give an expected value of $0.9490$, the observed maximum.
(a) One bootstrap sample from the cropped distribution.
(b) One realisation of the bootstrap sample in (a).
Figure 12: Histogram of cropped competition observations. Here the state of the art is an AUC of 0.8975 which leads to the expected observed maximum AUC in a competition being 0.9490.
Figure 12(a) shows one bootstrap sample from the kaggle observations cropped at $0.8975$. Number of classifiers is set to $m=3,173$, which is as many as there are classifiers in the melanoma challenge with AUC $>0.5$. Figure 12(b) shows a corresponding realisation, where the expected maximum is $0.9490$.
There are $2022$ teams with performance above the new SOTA candidate.
@@ 5.2.2 Correlation between classifiers in the AUC simulations
and equivalently for the negative class. However, the joint distribution of $(S_{0},S_{j})$ is not uniquely defined by the correlation, and simulations are not straightforward.
@@ 6 Discussion
We have addressed the mechanisms of multiplicity and how they affect the bias of the SOTA estimator. When accuracy is used as a performance measure, the exact probability distribution can be provided, and known analyses methods can be engaged. The bias in SOTA is substantial in our example, and possibly hinders new methods to attract interest from the scientific community. We have also demonstrated that dependency might not have a big impact on the bias when the probability of correct prediction is already high. The exception is when the dependency is very strong.
@@
This work shows that a multiplicity-adjusted state-of-the-art estimate can be calculated fairly easily through bootstrapping, and can serve as a better reference point when introducing new methods to the scientific community.

### v5

@@ 5.1.2 Cassava Leaf Disease
In this example, the maximum sample estimate is unlikely to be the product of multiplicity. For The Cassava Leaf Disease Classification competition, the task was to classify images of cassava leaves. The training set consisted of $21,367$ labelled images of five classes, and the test set of approximately $15,000$ images. $3,752$ teams participated (excluding those with sample performance lower than chance).
The three top performances appear as outliers (0.9132, 0.9043 and 0.9028) as seen in Figure 7(a). Several explanations for the outlying performances of the three top teams have been put forward, and the interested reader can take a deep-dive in the comment section of the competition. Since this is not the topic of the present work, we will not engage in the discussion here. The 95% CI of $\hat{\theta}=0.9132$ is $(0.9086,0.9177)$, no other sample performance is within this interval. By performing simulations without cropping, we get $E(\max(\hat{\theta}^{\prime}))=\max(\hat{\Theta})$, and although the upper tail is stretched, its tip does not surpass $\max(\hat{\Theta})$.
@@
Figure 7: Cassava Leaf sample performances and one realisation.
This analysis demonstrates that the suggested approach for SOTA estimation is robust in the sense that when multiplicity is not a likely explanation to the sample maximum value, the SOTA value is not altered by this approach.
5.2 Estimating $AUC_{SOTA}$
5.2.1 Melanoma Classification
@@
We will use the Melanoma Classification competition from 2020 as an illustrative example of $AUC_{SOTA}$ estimation because it follows many of the common structures of public challenges. The results are presented in Section 5.2.3.
@@
The SIIM-ISIC Melanoma Classification competition consisted of more than $40,000$ images of skin lesions from various sites, split into training ($33,126$), validation ($3,295$) and test set ($7,687$). The data set is highly imbalanced, with less than $2\%$ melanomas, which reflects the reality in many cancer detection situations. A full description can be found in Rotemberg et al. (2021). The size of the test set is smaller than the recommendations of, e.g., Roelofs et al. (2019), but reflects the reality of benchmark datasets and public challenges, as seen in, e.g., the overview of Willemink et al. (2020). We want to emphasise that the shortcomings of this competition are not unique, and that is precisely why it serves as a useful example.
Participants were allowed to use additional data for training if the data was shared on the competition web site. The prize money was set to $10,000$ USD for the best-performing method, measured as AUC on the hold-out test set. Figure 8 shows the results. Participants could get intermediate results from the validation set, published on a public leaderboard.
Figure 8: AUCs for the Melanoma Classification competition, calculated on a hold-out test set, cropped from below for better visualisation.
@@
The team ‘All Data Are Ext’, consisting of three Kaggle Grandmasters, won the competition, achieving an AUC of $0.9490$ on the test set, and their method and strategy is explained in Ha et al. (2020). Although we show that the AUC is an optimistic estimate for the melanoma detection SOTA, we do not question their rank in the challenge. Quite the contrary; the winning team chose many of the verified strategies from Section 2. They used additional data from previous years’ competitions, and they also did data augmentation. They chose ensemble learning as a strategy to avoid overfitting, a strategy implicitly supported by Talagrand (1996). They also trained on multiple classes, a strategy that Feldman et al. (2019) showed will slow down the overfitting.
In the following sections, we introduce a strategy for estimating $AUC_{SOTA}$, using simulated examples and the Melanoma Challenge for illustration.
5.2.2 Simulation of AUC
@@
In a binary classification problem, the calculation of AUC is based on scores of random variables from the positive and the negative class. The AUC is mathematically equivalent to the probability that a random variable from the positive class will be ranked higher, in terms of score, than a random variable from the positive class (Hanley and McNeil, 1982).
@@
Let $score(\cdot)$ be a function that assigns a score to a random variable. Let $S_{+}=score(W_{+})$, where $W_{+}$ is a random variable from the positive class, and $S_{-}=score(W_{-})$, where $W_{-}$ is a random variable from the negative class. Then $\text{AUC}=P(S_{-}<S_{+})=P(S_{-}-S_{+}<0)$, which is equivalent to the cdf at value 00. To simulate classifier scores according to a given AUC $=a$, we need to specify distributions for the variables $S_{-},S_{+}$ in such a way that $P(S_{-}-S_{+}<0)=a$. We use Normal distributions for illustration.
Let $S_{+}$ and $S_{-}$ have the distributions
@@
where $\mathcal{N}(\mu,\sigma^{2})$ is the Normal distribution with mean $\mu$ and variance $\sigma^{2}$. We then have that
@@
where $\Phi$ is the cdf of the standard Normal distribution. Specifying any three of the four parameters above decides the last one for a chosen AUC $=a$. Figure 9(a) shows an example where AUC = 0.90.
@@
Figure 9(a) shows an example of Normally distributed negative and positive scores and an AUC of 0.90. The rug shows an example draw with $n=3000$ and $\pi=0.017$, corresponding to the class balance in the Melanoma challenge.
(a) Score pdfs for a classifier with AUC= $0.90$; grey for negatives, black for positives. The rug shows an example draw from these with class balance $\pi=0.017$.
(b) ROC curves for a classifier with AUC $=0.90$ across 100 realisations of (a).
Figure 9: Example of what an AUC of $0.90$ might look like in practice. The right figure shows $100$ realizations of a ROC curves. The left figure shows the underlying score distributions that generated the curves in (b).
The choice of distributions has an impact on the final SOTA estimation, as it affects the variance of the ROC curves. However, with high AUC, the shape has less impact, and the analyst’s choices are not suggestive of the overall conclusion.
@@ 5.2.3 Simulation results for uncorrelated classifiers
Whereas the distribution of the sample maximum for probability of success is independent of class balance as long as the probability of success is independent of class, this is not true for the sample maximum of AUC, where the variance increases with class imbalance. We have chosen the same class balance as in the Melanoma challenge, $\pi=0.017$, as this reflects real world examples. For simplicity we choose $\mu_{-}=0$ and $\sigma_{-}=\sigma_{+}=1$. We then solve for $\mu_{+}$ numerically. Now we can generate scores from the resulting Normal distribution. The 95% CI of a single classifier with $\hat{AUC}=0.90$ is (.$(.8558,0.9376)$, and gives a strong indication of what we can expect when $1,000$ teams apply their classifiers on a test set.
Figure 10: Simulated distribution of the sample maximum AUC with parameters as in Figure 9. This is an analogous figure to Figure 2(b) in Section 3.4.
Figure 10 shows the simulated distribution of the sample maximum AUC over $10,000$ repetitions of a competition where every classifier has a true AUC of $0.90$. The simulated expected value is $0.9562$, the standard deviation is $0.004459$, and the simulated 95% CI is $(0.9486,0.9662)$. Comparing this to the results in Section 3.4, we see that the bias is much larger for the sample maximum AUC. This is partly due to the large variance created by the class imbalance. Although the sample size is $n=3,000$, only $51$ of those belong to the positive class, creating large variance in the estimator.
Again, it is clear that the observed maximum is an unlikely SOTA candidate, and to give a better estimate of the SOTA for the Melanoma challenge, we follow the same procedure as in Section 5.1: The competition sample estimates are cropped in such a manner that their realisations give an expected value of $0.9490$, the competition sample maximum.
(a) Truncated AUC distribution based on the Melanoma challenge scores. The part of the distribution removed by truncation is shaded in dark grey.
(b) A realization from the truncated distribution of (a).
Figure 11: Histogram of cropped competition AUC sample estimates. Here the state of the art is an AUC of 0.8975 which leads to the expected sample maximum AUC of 0.9490.
Figure 11(a) shows one bootstrap sample from the kaggle observations cropped at $0.8975$. Number of classifiers is set to $m=3,173$, which is as many as there are classifiers in the melanoma challenge with AUC $>0.5$. Figure 11(b) shows a corresponding realisation, where the maximum is close to $0.9490$. There are $2,022$ teams with performance above the new SOTA candidate.
The discrepancy between the sample maximum and the new SOTA candidate is big, and have big implications. From an AUC close to $95\%$ to just under $90\%$ is a substantial drop in performance in a cancer detection setting. In addition, two-thirds of the classifiers have performances above the new SOTA estimate. As was shown in Section 4, illustrated in Figure 4(b), the bias is reduced for correlated classifiers, so the analysis with independence assumption is an overestimation of the bias. In this challenge, we can assume multiple sources of dependency: Access to the same training set, models pre-trained on the same data, similar models, etc. However, the wide CI for a single classifier tells the story of an estimator with large variance, and hence, the bias of the sample maximum will be substantial.
@@ 5.2.4 Correlation between classifiers in the AUC simulations
and equivalently for the negative class. However, the joint distribution of $(S_{0},S_{j})$ is not uniquely defined by the correlation, and additional assumptions must therefore be made. We leave this as future work.
@@ 6 Discussion
We have addressed the mechanisms of multiplicity and how they affect the bias of the SOTA estimator. When probability of success is used as a performance measure, the exact distribution of the sample maximum estimator can be provided in the simpler cases, and known analyses methods can be engaged. The bias in SOTA is substantial in our examples, and possibly hinders new methods to attract interest from the scientific community.
@@
This work shows that for some cases, an unbiased SOTA estimate can be calculated fairly easily through bootstrapping, and can serve as a better reference point when introducing new methods to the scientific community. In other cases, the SOTA estimation is more complicated.